# Bringing Characters to New Stories: Training-Free Theme-Specific Image Generation via Dynamic Visual Prompting

### v1

@@ 6. Limitations
Figure 13. Results of user-specific refinement.
Figure 14. Failure cases due to data limitation.
T-Prompter can generate diverse actions and backgrounds that do not exist in the dataset, but it may be limited by overly homogeneous or insufficient data. We recommend that users provide images of the target characters in multiple poses and varied backgrounds. As shown in Fig. 14, if only facial images of a character are provided and the model is tasked with generating a full-body image, the results may appear reasonable but lack consistency with the actual target. Beyond collecting richer data, this limitation could be addressed through data augmentation leveraging generative models.
@@ 7. Conclusion
This paper proposes T-Prompter, a novel training-free, modification-free method of theme-specific image generation with high flexibility and accuracy. We introduce visual prompting, a form of interaction with generative models, which can provide more accurate and direct guidance for models in the visual domain. Our dynamic visual prompting pipeline leverages the capabilities of multi-modal models and LLMs while utilizing data-driven advantages to meet the demanding requirements of flexible theme-specific generation. In comparisons with several state-of-the-art baseline methods, T-Prompter achieves superior results in both qualitative and quantitative evaluations. Following a training-free technical paradigm, T-Prompter delivers performance comparable to fine-tuned models, with the added benefit of producing highly realistic outputs. Experimental results demonstrate the feasibility and effectiveness of T-Prompter across a variety of applications, and its low cost lowers the barrier to use. Additionally, the DVP working mode highlights the potential of inpainting-based visual models in enhancing image generation processes. We believe that more targeted training and fine-tuning approaches will further unlock the potential of more efficient, controllable generation in future applications.
Figure 15. Results of story generation.
Figure 16. Results of realistic character generation.
Figure 17. More results of style-guided image generation.
Figure 18. More results of story generation.
Figure 19. More results of new character generation.

### v2

@@ 6. Limitations
IP-Prompter can generate diverse actions and backgrounds that do not exist in the dataset, but it may be limited by overly homogeneous or insufficient data. We recommend that users provide images of the target characters in multiple poses and varied backgrounds. As shown in Fig. 12, if only facial images of a character are provided and the model is tasked with generating a full-body image, the results may appear reasonable but lack consistency with the actual target. Beyond collecting richer data, this limitation could be addressed through data augmentation leveraging generative models.
@@ 7. Conclusion
This paper proposes IP-Prompter, a novel training-free, modification-free method of theme-specific image generation with high flexibility and accuracy. We introduce visual prompting, a form of interaction with generative models, which can provide more accurate and direct guidance for models in the visual domain. Our dynamic visual prompting pipeline leverages the capabilities of multi-modal models and LLMs while utilizing data-driven advantages to meet the demanding requirements of flexible theme-specific generation. In comparisons with several state-of-the-art baseline methods, IP-Prompter achieves superior results in both qualitative and quantitative evaluations. Following a training-free technical paradigm, IP-Prompter delivers performance comparable to fine-tuned models, with the added benefit of producing highly realistic outputs. Experimental results demonstrate the feasibility and effectiveness of IP-Prompter across a variety of applications, and its low cost lowers the barrier to use. Additionally, the DVP working mode highlights the potential of inpainting-based visual models in enhancing image generation processes. We believe that more targeted training and fine-tuning approaches will further unlock the potential of more efficient, controllable generation in future applications.
Acknowledgements.
This work was supported in part by the Beijing Natural Science Foundation under nos. L221013 and QY24384, in part by the Beijing Training Program of Innovation and Entrepreneurship for Undergraduates under no. S202414430024, in part by the National Science and Technology Council under no. 111-2221-E-006-112-MY3, Taiwan, in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy-EXC 2117-422037984.